OpenStudy (anonymous):

f(x) = sin (x/2) - cos x find the zeros using half angle formulas

OpenStudy (anonymous):

\[ \sin(x/2)=\pm\sqrt{\frac{1-\cos x}{2}} \]

OpenStudy (anonymous):

how should I approach that after plugging that in?

OpenStudy (anonymous):

Set it equal to \(\cos x\) and square both sides.

OpenStudy (anonymous):

thanks, brain isnt working today